# Revisiting Multi-Agent Asynchronous Online Optimization with Delays: the Strongly Convex Case

### v1

@@ 3.3 Approximate Follow-the-Delayed-Leader
Figure 1: Comparisons of our A-FTDL against DDA on delayed $2$-agent online binary classification.
@@ 4 Experiments
Dataset # Examples ($T$) # Features ($n$) # Classes
@@
Specifically, we consider a delayed $2$-agent online binary classification problem with strongly convex loss functions. For each dataset, we utilize all training examples to conduct our experiments, and thus $T$ is equal to the number of these training examples. In each round $t\in[T]$, one of the agent will be active, and choose a decision $\mathbf{x}_{t}$ from the following convex set
@@
which satisfies Assumption  2 with $D=2$, where $n$ is equal to the number of features in the dataset. Then, an example $(\mathbf{w}_{t},y_{t})$ is selected from the dataset, and causes a loss
@@
Note that by a single pass of the datasets, we can compute the maximum norm of $\mathbf{w}_{t}$, and denote this value as $w_{max}$. Then, the above loss function $f_{t}(\mathbf{x})$ also satisfies Assumption 1 with
$G=\frac{\lambda D}{2}+w_{max}=0.01+w_{max}.$
Next, to simulate the delays, we randomly generate two delayed sequences for the two agents, each sequence is divided into four parts with delays selected from ranges of $[1,250]$, $[1,500]$, $[1,750]$, and $[1,1000]$, respectively. In this way, the maximum delay $d$ is equal to $1000$. Moreover, we recall that the DDA algorithm needs to choose an appropriate $\mathcal{R}(\mathbf{x})$ and tune the parameter $\eta_{t}$. According to Proposition 7 of Hsieh et al. (2022), we set $\mathcal{R}(\mathbf{x})=(1/2)\|\mathbf{x}\|_{2}^{2}$ and use
$\eta_{t}=\frac{D}{2\sqrt{2}G\sqrt{(1+2d)(|\mathcal{F}_{t}|+d+1)}}$
@@
Fig. 1 shows the comparisons of our A-FTDL and the DDA algorithm (Hsieh et al., 2022) over all these datasets. We find that the cumulative loss of A-FTDL is smaller than that of DDA, which verifies the theoretical improvement on the regret by exploiting the strong convexity of functions. Moreover, these results also demonstrate the benefit of strong convexity for removing the requirement on prior information and assumption about the delay.
@@ 5 Conclusion and Future Work
In this paper, we investigate multi-agent asynchronous online optimization with delays, and develop two novel algorithms called FTDL and A-FTDL for the case of strongly convex functions. According to the analysis, our two algorithms exploit the strong convexity to achieve an $O(d\log T)$ regret bound, which is much tighter than the existing $(\sqrt{dT})$ regret bound achieved by only using the convexity. More interestingly, different from the existing algorithm that needs to know the maximum delay beforehand or requires an additional assumption on the arrival order of feedback, our two algorithms do not require any prior information or assumption about the delay. Finally, numerical experiments show that our A-FTDL outperforms the existing algorithm in real applications.
An open question is whether our algorithms and analysis can be further extended to a broader class of exponentially concave (exp-concave) functions. We notice that in the standard OCO, there is also a variant of FTL (Hazan et al., 2007) that enjoys an $O(n\log T)$ regret bound for exp-concave functions. However, it seems highly non-trivial to exploit the exponential concavity in the delayed setting, even with only single agent (Wan et al., 2022b). Therefore, we leave this extension for future work.

### v2

@@ 3.3 Approximate Follow-the-Delayed-Leader
Figure 1: Comparisons of our A-FTDL against DDA on delayed 22-agent online binary classification.
3.4 Extension to Unconstrained Domains
Furthermore, we consider the unconstrained setting where $\mathcal{K}=\mathbb{R}^{n}$, and extend our algorithm to handle this case while maintaining strong theoretical guarantees. Note that such an extension follows from existing studies on delayed OCO with strongly convex functions (Wan et al., 2024a). Here, we provide the details for completeness. To be precise, in unconstrained domains, the original Assumptions 1 and 2 may not hold globally. To address this challenge, we introduce two weaker and more reasonable assumptions.
  • All loss functions are $G$-Lipschitz continuous at the origin, i.e., $|f_{t}(\mathbf{0})-f_{t}(\mathbf{x})|\leq G\|\mathbf{x}\|_{2}$ for all $\mathbf{x}$ in a neighborhood of $\mathbf{0}$.
  • Each loss function $f_{t}$ is $\lambda$-strongly convex over $\mathbb{R}^{n}$.
Under these two weaker assumptions, by applying Lemma 2, it is not hard to verify that,
$\|\mathbf{x}^{*}\|_{2}\leq\frac{2}{\lambda}\left(f_{t}(\mathbf{0})-f_{t}(\mathbf{x}^{*})\right)\leq\frac{2G}{\lambda}$ (15)
where $\mathbf{x}^{*}=\operatorname*{argmin}_{\mathbf{x}\in\mathbb{R}^{n}}\sum_{t=1}^{T}f_{t}(\mathbf{x})$. As a consequence, the player only needs to select actions from the constrained set:
$\mathcal{K}^{\prime}=\left\{\mathbf{x}\in\mathbb{R}^{n}\middle|\|\mathbf{x}\|_{2}\leq\frac{2G}{\lambda}\right\}$ (16)
which naturally satisfies Assumption 2 with radius $R=2G/\lambda$. We further assume that all loss functions satisfy the original Assumptions 1 and 3 over the restricted set $\mathcal{K}^{\prime}$. This allows us to reduce the unconstrained problem to a constrained one over $\mathcal{K}^{\prime}$. Theorem 1 shows that the regret bound of FTDL is independent of the domain’s radius and therefore remains valid. For A-FTDL, we derive the regret bound from Theorem 2 by applying Algorithm 2 over a bounded domain of radius $\frac{2G}{\lambda}$, then we have
$R_{T}\leq\frac{50dG^{2}(\ln T+1)}{\lambda}.$
@@ 4 Experiments
Dataset # Examples # Features (i.e., $n$) # Classes
@@
Specifically, we randomly select $10000$ examples from the original data set. Then we use the first $8000$ data points as the training set and the subsequent $2000$ data as the test set. Thus, the total number of rounds $T$ in our experiments is $8000$.
We consider a delayed multi-agent online binary classification problem with strongly convex loss functions. Under two scenarios: (i) a system with $2$ agents and a maximum delay of $100$, and (ii) a system with $20$ agents and a larger maximum delay of $1000$. These settings are inspired by the practical observation that as the number of agents increases, the communication and coordination overhead often grows, leading to longer feedback delays. In each round $t\in[T]$, one of the agents will be active, and choose a decision $\mathbf{x}_{t}$ from the following convex set
@@
which satisfies Assumption 2 with $R=1$, where $n$ is equal to the number of features in the dataset. Then, an example $(\mathbf{w}_{t},y_{t})$ is selected from the dataset, and causes a loss
@@
Note that by a single pass of the datasets, we can compute the maximum norm of $\mathbf{w}_{t}$, and denote this value as $w_{\max}$. Then, the above loss function $f_{t}(\mathbf{x})$ also satisfies Assumption 1 with
$G=\lambda R+w_{\max}=0.01+w_{\max}.$
Next, to simulate the delays, we independently generate random delay sequences for each agent. For the 2-agent scenario, each agent’s delays are randomly sampled from the range $[1,100]$. For the 20-agent scenario, each agent’s delays are randomly sampled from the range $[1,1000]$. Moreover, we recall that the DDA algorithm needs to choose an appropriate $\mathcal{R}(\mathbf{x})$ and tune the parameter $\eta_{t}$. According to Proposition 7 of Hsieh et al. (2022), we set $\mathcal{R}(\mathbf{x})=(1/2)\|\mathbf{x}\|_{2}^{2}$ and use
$\eta_{t}=\frac{R}{\sqrt{2}G\sqrt{(1+2d)(|\mathcal{F}_{t}|+d+1)}}$
@@
Fig. 1 shows the comparisons of our A-FTDL and the DDA algorithm (Hsieh et al., 2022) on the two settings. We observe that in both settings, the cumulative loss of A-FTDL is consistently lower than that of DDA, verifying the theoretical improvement on regret by exploiting strong convexity. Moreover, as the number of agents increases and the delays become larger, the performance gap between A-FTDL and DDA widens, further demonstrating the benefit of strong convexity for removing the need for prior information and assumptions about the delay.
Table 2 shows that the running times of A-FTDL and DDA are comparable under both settings. Particularly, in the 2-agent setting with smaller delays, the difference between the two algorithms is even smaller, which is consistent with our previous complexity analysis.
Finally, we evaluate the test accuracy of the final decision vector $\mathbf{x}_{T}$ obtained by each algorithm on the $2000$ test examples. The evaluation is conducted under the setting of 2-agent with $d_{\max}=100$ and 20-agent with $d_{\max}=1000$. From Table 3, our A-FTDL has higher testing accuracy than DDA in two datasets: a9a and w8a. In the other two datasets, A-FTDL and DDA have the same testing accuracy. We want to emphasize that these results are reasonable, because the testing accuracy actually is not directly determined by the regret bound.
Dataset 2 agents, $d=100$ 20 agents, $d=1000$
A-FTDL DDA A-FTDL DDA
a9a 0.1970 0.1465 1.6503 1.2387
ijcnn1 0.1338 0.1112 1.2645 1.1594
phishing 0.1271 0.1377 1.3161 1.2222
w8a 0.2271 0.1494 1.5526 1.1928
Table 2: Average per-round running time (seconds) comparison on different datasets
Dataset 2 agents, $d=100$ 20 agents, $d=1000$
A-FTDL DDA A-FTDL DDA
a9a 83.70% 76.05% 82.55% 76.00%
ijcnn1 90.05% 90.05% 89.80% 89.80%
w8a 89.80% 89.65% 89.80% 89.60%
phishing 56.60% 56.60% 56.40% 56.40%
Table 3: Test accuracy (%) comparison on on different datasets
@@ 5 Conclusion and Future Work
In this paper, we investigate multi-agent asynchronous online optimization with delays, and develop two novel algorithms called FTDL and A-FTDL for the case of strongly convex functions. According to the analysis, our two algorithms exploit the strong convexity to achieve an $O(d\log T)$ regret bound, which is much tighter than the existing $O(\sqrt{dT})$ regret bound achieved by only using the convexity. More interestingly, different from the existing algorithm that needs to know the maximum delay beforehand or requires an additional assumption on the arrival order of feedback, our two algorithms do not require any prior information or assumption about the delay. Finally, numerical experiments show that our A-FTDL outperforms the existing algorithm in real applications.
An open question is whether our algorithms and analysis can be further extended to a broader class of exponentially concave (exp-concave) functions. We notice that in the standard OCO, there is also a variant of FTL (Hazan et al., 2007) that enjoys an $O(n\log T)$ regret bound for exp-concave functions. Very recently, Qiu et al. (2025) have proposed algorithms for delayed OCO with exp-concave functions, achieving a regret bound of $O(dn\log T)$. However, the extension to multi-agent asynchronous online optimization seems still highly non-trivial. Therefore, we leave this extension for future work.
Acknowledgements
This work was partially supported by National Natural Science Foundation of China (62306275) and Yongjiang Talent Introduction Programme (2023A-193-G). The authors would also like to thank the anonymous reviewers for their helpful comments.